OpenStudy (chillhill):

The quadratic equation 4x^2 + 4kx + 9 = 0, k > 0 has exactly one solution for x. Find the value of k.

OpenStudy (michele_laino):

hint: we have to require that the discriminant of such trinomial, is zero: \[\Large \Delta = {\left( {4k} \right)^2} - \left( {4 \cdot 36} \right) = 0\] and then we have to consider the positive solution only of such equation for \(k\)